OpenStudy (anonymous):

Franco bought 1.5 pounds of deli cheese for$6.75 . If he went back for another 2 pounds of the same cheese, how much could Franco expect to pay?

OpenStudy (anonymous):

set up a proportion... 1.5/6.75 = 2/x (1.5 lbs for 6.75 = 2 lbs for x$) now cross multiply... (1.5)(x) = (6.75)(2) now solve for x can you take it from here ?

OpenStudy (destinymasha):

9.00

OpenStudy (anonymous):

8 dolers and 44 cents

OpenStudy (anonymous):

1.5x = 13.50 x = 13.50/1.5 x = 9.00 Destiny, you got it right . Good job :)